Applied Cryptography Assignment

Using the extended Euclidean algorithm, find the multiplicative inverse of:

13140 mod 40902
Answer #1

Answer:

Given that 13140 mod 40902.

Multiplicative inverse:

=gcd(40902,13140)

=gcd(13140,1482)

=gcd(1482,1284)

=gcd(1284,198)

=gcd(198.96)

=gcd(96,6)

=gcd(6,0)

So, the GCD or HCF (40902, 13140)

6 1.

So, there is no multiplicative inverse .....

>>>>>>>> Alright Dude . I hope this explanation will help you . if it works hit THUMSUP or incase if you have any doubts .... just drop your comment ...THANK YOU<<<<<<<<<<<<<<